OpenStudy (anonymous):

can someone help me with this. Add the fraction 5/7+3/8 and write the result in simplest form

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 5/7+3/8\LARGE\color{white}{ \rm │ }}\). can you first find the common denominator ?

OpenStudy (anonymous):

56?

OpenStudy (solomonzelman):

YEs !

OpenStudy (solomonzelman):

Okay, so you will get \(\Large\color{blue}{ \frac{5}{7}+\frac{3}{8}\LARGE\color{white}{ \rm │ }}\) \(\Large\color{blue}{ \frac{5\times 8}{56}+\frac{3\times 7}{56}\LARGE\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

can you multiply on top of each fraction, and add the factions together ?

OpenStudy (solomonzelman):

Add the *fractions

OpenStudy (anonymous):

ok is that called cross multiplying

OpenStudy (solomonzelman):

no, cross multiplying, is t=when you have 2 fractions equal to each other, and in one of the fractions you have an unknown variable, and you then cross multiply.

OpenStudy (solomonzelman):

you JUST multiply

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \frac{5\times 8}{56}+\frac{3\times 7}{56}\LARGE\color{white}{ \rm │ }}\) and I will get, \(\Large\color{blue}{ \frac{40}{56}+\frac{21}{56}\LARGE\color{white}{ \rm │ }}\) right ?

OpenStudy (anonymous):

right

OpenStudy (solomonzelman):

add the fractions (without adding the denominators, which are the bottoms of the fractions) and tell me what you get.

OpenStudy (anonymous):

61/56

OpenStudy (solomonzelman):

Yes.

OpenStudy (solomonzelman):

You can either write it as an improper fraction (the way you have it right now) OR, you can say that it is 1 5/56. Which is te same exact thing, but in terms of a mixed number.

OpenStudy (solomonzelman):

Good work :)

OpenStudy (anonymous):

thank you so much! but how do you write it in simplest form?

OpenStudy (solomonzelman):

It would be either, \(\large\color{blue}{ \frac{61}{56}\LARGE\color{white}{ \rm │ }}\), or \(\large\color{blue}{ 1\frac{5}{56}\LARGE\color{white}{ \rm │ }}\).

OpenStudy (solomonzelman):

I would use the mixed number though, since it is actually properly written.

OpenStudy (anonymous):

oh ok thanks again!

OpenStudy (solomonzelman):

Anytime !